How to solve these problems? 4w-3w+2w=24
4(x+3)=20
-8+t/3=7
4y+3
____=9
7



Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

4W-3W+2W=24
3W=24
W=24/3
W=8 PROOF 4*8-3*8+2*8=24
32-24+16=24
48-24=24
24=24
----------------------------------------------------
4(X+3)=20
4X+12=20
4X=20-12
4X=8
X=8/4
X=2
-----------------------------------------------------
-8+T/3=7
T/3=7+8
T/3=15
T=15*3
T=45
-------------------------------------------------------
(4Y+3)/7=9
4Y+3=9*7
4Y+3=63
4Y=63-3
4Y=60
Y=60/4
Y=15

Answer by tutorcecilia(2152)


4w-3w+2w=24
1w+2w=24
3w=24
3w/3=24/3
w=8
Check by plugging (w=8) back into the original equation.
.
4(x+3)=20
4x+12=20
4x+12-12=20-12
4x=8
4x/4=8/4
x=2
Check by plugging (x=2) back into the original equation.
.
-8+t/3=7
-8-8+t/3=7+8
t/3=15
(3)(t/3)=(15)(3)
t=45
Check by plugging (t=45) back into the original equation]
.
(y+3)/7=9
(7)(y+3)=(9)(7)
y+3=63
y+3-3=63-3
y=60
Check by plugging (y=60) back into the original equation.